### v2

# An Ecosystem for Ontology Interoperability

Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
Index Terms:
ontology, data interoperability
I Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a logic-based expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain. There is no guarantee that ontologies will agree on the same conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies. For example, it is common for ontologies to encompass a range of different concepts. A person can be described using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee). Moreover, the terminology used in ontologies can be different. It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
II Related Work
II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2]. Fig. 1 illustrates that ODPs use the star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
Figure 1: Ontology design patterns (ODPs).
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies. They are static and are not prone to capturing changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
II-B Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. Fig. 2 shows that OM&OV conform to the pairwise matching schema. This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
Figure 2: Ontology matching and versioning (OM&OV).
OM&OV are typically resource-consuming. The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
II-C Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) [5]. Fig. 3 shows that the OCKGs comply with the star-based matching schema. While ontologies comply with KG(s), their reference mappings can be derived from the concepts pointing to the same data instance.
Figure 3: Ontology-compliant knowledge graphs (OCKGs).
Building OCKGs is not easy. This process requires extensive workloads to find the alignments between different ontologies, so that each KG can be modelled with different ontologies to determine which ontology fits the KG(s) better.
III An Ecosystem
Fig. 4 illustrates an overview of the ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
Figure 4: An overview of the ecosystem.
The design phase is responsible for understanding the requirements and creating a blueprint or prototype. In ontology engineering, these requirements are collected by competency questions. ODPs can serve as a skeleton to create conceptual models based on competency questions. The develop phase includes turning the prototype into a full product and applying a unit test. In ontology engineering, this phase is referred to as ontology construction and evaluation. OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version. The deploy phase usually refers to implementing the product in the real world. In ontology engineering, we refer to this as applying the ontology to actual KGs. We can utilise OCKGs to comprehend the interoperability of ontologies within each KG or a group of KGs, thereby enhancing the design of the ontology in the next cycle.
IV A Case Study in the Building Domain
Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades. While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain [6]. We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability. The running example is described as follows: A sensor is installed in the room to monitor the room temperature. The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is recommended to build an ontology based on ODPs. Several ODPs can be used in this case, as illustrated in Fig. 5.
  1. (1)
    Observation Pattern (used in SOSA [7] and SSN [8]). Sensor observation is defined as a type of observation. The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
  2. (2)
    Measurement Pattern (used in the early version of SAREF [9]). Sensor observation is defined as a measurement. The measurement is made by the sensor, which is a measure of the room related to the property temperature.
  3. (3)
    Event Pattern (used in DUL [10]). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
  4. (4)
    Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
While all of these ODPs can effectively capture the main concepts in the example, they have differences that need to be considered.
  1. 1.
    The definition of the concept can be different. For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
  2. 2.
    The levels of expressiveness are different. Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish whether it is a structural attachment (i.e. isPartOf) or only a spatial presence (i.e. isLocatedIn).
Figure 5: Modelling using different ODPs.
IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN. As illustrated in Fig. 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. If we consider SOSA to be an updated version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [12]. OM&OV tasks can be classified into two main categories.
Figure 6: Aligning SOSA and SSN.
  1. 1.
    Finding trivial alignments only considers the syntactic matching of the entity. For example, the mappings between the classes sosa:Sensor and ssn:Sensor, and the properties sosa:madeBySensor and ssn:madeBySensor. We also refer to these mappings as “seed mappings” because they serve as the basis for subsequent lexical and semantic matching.
  2. 2.
    Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity. For example, the mappings between the classes sosa:ObservableProperty and ssn:Property, and the properties sosa:isHostedBy and ssn:onPlatform. The alignment should also exclude non-existent and counter-intuitive mappings. For example, ssn:hasProperty should not have a direct equivalent mapping in SOSA.
IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it remains challenging to determine which ontology to use in the real world. Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
  1. 1.
    The stimulus of the sensor “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
  2. 2.
    The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings (retrieved from [13, 14, 15]).
  1. 1.
    We use ssn:Property to replace sosa:ObservableProperty because this class is more general and can be applied in a broader context for sensing, including actuation and sampling.
  2. 2.
    We use sosa:isHostedby to replace ssn:onPlatform because this property is more closely related to describing a sensor physically installed in the room.
  3. 3.
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
Figure 7: Build OCKGs using SOSA and SSN.
Fig. 8 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is more suitable for the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
Figure 8: Updated ontology and new OCKG.
IV-D Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align the concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
V Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Introduced in 2023, ontology-compliant KGs aim to explore the interoperability between ontology and its complementary KGs.
While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and constructing OCKGs in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
Supplemental Material Statement
The source code, data, and/or other artifacts used in this study have been made available at https://github.com/qzc438-research/ontology-interoperability.
Acknowledgment
The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.