### v6

# ISWC 2026 Companion Volume, October 25-29, 2026, Bari, Italy

Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and data-driven ontology validation (DOVA) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
keywords:
[…]
1 Problem Statement
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability among KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a class schema expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a person, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application. There is no guarantee that ontologies in the same domain share the same conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects the sharing and reuse of ontologies. Ontologies can encompass different entities to describe the same concepts. For example, describing a person can use their basic profiles (name, age, and gender), social relationships (parent-child and friendship), or professional roles (student and employee). The terminology used in the ontologies may also vary. One ontology can use “first name” and “last name” to describe a person’s full name, and another ontology can use “given name” and “family name” instead.
To tackle this challenge, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA). We analyse their pros and cons and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to demonstrate how these three techniques can foster strengths and circumvent weaknesses throughout the ontology engineering life cycle.
2 Related Work
2.1 Ontology Design Patterns (ODPs)
Ontology design patterns (ODPs) identify common ontology fragments across different ontologies [2]. ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies (see Figure 1). ODPs have some limitations. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control terminology, and ontology creators remain free to make their own choices. Secondly, ODPs are commonly handcrafted based on existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
2.2 Ontology Matching and Versioning (OM&OV)
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. OM&OV conform to a pairwise matching schema. This means that each mapping is a pairwise process to find the correspondences between two ontologies (see Figure 1). OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
2.3 Data-Driven Ontology Validation (DOVA)
Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology. DOVA complies with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties that point to the same KG entities (see Figure 1). The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Building OCKGs is not easy. To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate ontology. This process essentially involves aligning the candidate ontologies, informed by the KG under study.
Figure 1: Illustration of different matching schema used in ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA).
3 An Ecosystem for Ontology Interoperability
Figure 2 illustrates an overview of the ecosystem for ontology interoperability. Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ODPs, OM&OV, and DOVA in different phases. (1) The design phase is responsible for understanding the requirements and creating a blueprint or prototype. In ontology engineering, these requirements are collected by competency questions. ODPs can serve as a skeleton to create conceptual models based on competency questions. (2) The develop phase includes turning the prototype into a full product and applying a unit test. In ontology engineering, this phase is referred to as ontology construction and evaluation. OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version. (3) The deploy phase usually refers to implementing the product in the real world. In ontology engineering, we refer to this as applying the ontology to actual KGs. We can utilise OCKGs to assess the interoperability of ontologies within a single KG or across a set of KGs, thereby enhancing the design of the ontology in the next cycle.
Figure 2: An overview of the ecosystem. We propose to use ODPs in the design phase, OM&OV in the develop phase, and DOVA in the deploy phase to achieve better ontology interoperability.
4 A Case Study in the Building Domain
Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades. While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain [6]. We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability. The running example is described as follows: A sensor is installed in the room to monitor the room temperature. The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
4.1 Ontology Design Patterns (ODPs) in the Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Figure 3 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
  1. (i)
    Observation Pattern (used in SOSA [7] and SSN [8]). Sensor observation is defined as a type of observation. The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
  2. (ii)
    Measurement Pattern (used in the early version of SAREF [9]). Sensor observation is defined as a measurement. The measurement is made by the sensor, which is a measure of the room related to the property temperature.
  3. (iii)
    Event Pattern (used in DUL [10]). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
  4. (iv)
    Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
  1. (i)
    The definitions of the concepts can differ. Temperature is defined as a property of the room in both the Observation Pattern and the Measurement Pattern, but in the Event Pattern it is treated as a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
  2. (ii)
    The levels of expressiveness are different. Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish between a structural attachment (i.e. isPartOf) and a spatial presence (i.e. isLocatedIn).
Figure 3: Modelling sensor observation using different ODPs.
4.2 Ontology Matching and Versioning (OM&OV) in the Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Figure 4, SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to identify aligned concepts across two ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
Figure 4: Aligning SOSA and SSN.
  1. (i)
    Finding trivial alignments only considers the syntactic matching of the entity. For example, the mappings between the classes sosa:Sensor and ssn:Sensor, and the properties sosa:madeBySensor and ssn:madeBySensor. We also refer to these mappings as “seed mappings” because they serve as the basis for subsequent lexical and semantic matching.
  2. (ii)
    Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity. For example, the mappings between the classes sosa:ObservableProperty and ssn:Property, and the properties sosa:isHostedBy and ssn:onPlatform. The alignment should also exclude non-existent and counter-intuitive mappings. For example, ssn:hasProperty should not have a direct equivalent mapping in SOSA.
4.3 Data-Driven Ontology Validation (DOVA) in the Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Figure 5 shows the OCKGs based on SOSA and SSN. We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
Figure 5: Build OCKGs using SOSA and SSN.
  1. (i)
    The stimulus of the sensor “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
  2. (ii)
    The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA, as well as make several design choices between aligned mappings (retrieved from [13, 15, 16]).
  1. (i)
    We use ssn:Property to replace sosa:ObservableProperty because this class is more general and can be applied in a broader context for sensing, including actuation and sampling.
  2. (ii)
    We use sosa:isHostedby to replace ssn:onPlatform because this property is more closely related to describing a sensor physically installed in the room.
  3. (iii)
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we add the verb “has” to distinguish the property from the class.
Figure 6 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is better suited to the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
Figure 6: Updated ontology and new OCKG.
4.4 Reflection and Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
5 Contributions Made So Far
  1. (i)
    ODPs: We systematically analyse four popular ontologies used in the smart building domain and observe that each building ontology has merits and demerits. We propose building ODPs to capture common and well-agreed concepts and relations in the building domain, including building spaces, equipment and systems, points and measurements [6].
  2. (ii)
    OM&OV: We study the interlink between modern LLMs and traditional OM systems [17]. We develop a novel software system for OM and OV. Agent-OM [18] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner. Agent-OV [14] evolves from Agent-OM and explicitly targets OV tasks, including new task formulation and evaluation, with a cross-reference mechanism that compiles OM tasks. Agent-OM is used to create an LLM hallucination dataset [19] for the LLM leaderboard and LLM fine-tuning for OM. Agent-OM also participates in the Ontology Alignment Evaluation Initiative (OAEI) [20] 2025 campaign and showcases its generative capability in handling OM tasks at different levels of granularities [21].
  3. (iii)
    DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance. Examples in the building domain demonstrate the process of constraining OCKGs and showcase the efficiency of OCKGs in facilitating ontology validation and decision-making [5].
6 Conclusion and Future Work
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Data-driven ontology validation, which focuses on exploring the interoperability between an ontology and its complementary KGs, has received increasing attention in recent years. While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve ontology interoperability throughout the ontology engineering life cycle.
Acknowledgements.
The author is supervised by Kerry Taylor (Australian National University, ANU), Weiqing Wang (Monash University), Jing Jiang (ANU), Subbu Sethuvenkatraman (Commonwealth Scientific and Industrial Research Organisation, CSIRO), Qing Wang (ANU), and Alex Potanin (ANU). The author thanks CSIRO for supporting this project.
Declaration on Generative AI
During the preparation of this work, the authors used Grammarly in order to grammar and spell check, and to improve the text readability. After using the tool, the authors reviewed and edited the content and take full responsibility for the publication’s content.